# Magnetotransport Properties of Epitaxial Films and Hall Bar Devices of the Correlated Layered Ruthenate SrRuO

### 15 Dec 2023 v3

@@ 3. Results and Discussion
The crystalline quality and phase purity of SrRuO films were characterized by XRD. Figure 2(a) shows representative scans for the SrRuO films at various substrate temperatures and laser fluence. For the films grown at optimal growth conditions (TC and laser fluence ), only the substrate peaks and phase-pure family of the film diffraction peaks are resolved, which indicates a high crystallinity and verifies that the SrRuO films were aligned along the c-axis. The Bragg diffraction peaks are well resolved, while the peaks are not present due to the small film thickness and the low structure factors of these peaks [10]. The extracted out-of-plane lattice parameter for epitaxial films grown at the optimal growth conditions is Å. This value is consistent to the bulk axis lattice constant of 20.7194 Å reported previously in single crystals of SrRuO [40, 41, 42, 30]. Extra phases, indicated by in Fig. 2(a), were detected in SrRuO films grown at high substrate temperatures, underlining the role of proper tuning of T for achieving phase stability in these films. These extra phases are associated with the family of the diffraction peaks of the R-P phase [43, 44], in agreement with STEM data [Fig. S1].
Figure 2(b) shows a closeup view of the scans around the diffraction peak for the SrRuO films grown at various T and laser fluence. For the films grown at the optimal growth conditions, only the peaks are resolved with noticeable thickness fringes, which highlight phase purity and smooth growth. However, we observed shifting of the peaks to lower angles for films that were not grown at the optimal growth conditions. The shifting of the peak demonstrates that the -axis lattice constant expands, indicating inter-growth of other SrRuO phases and change of stability of the SrRuO films. In particular, the SrRuO films grown at T°C exhibited a peak shoulder around the substrate peak, indicating an inter-growth of higher R-P phases, with predominance of the SrRuO phase [46]. These observations are consistent with the STEM data [see, Fig. S1] in which considerable inter-growths of other members of the SrRuO series were clearly detected for epitaxial SrRuO films prepared outside the optimal growth conditions.
@@
For epitaxial SrRuO films grown at optimal growth conditions (TC and laser fluence ), the defects detected are infrequently regions of inter-growth within layers (indicated by black arrows in Figs. 3(a) and 3(c)). Such inter-growths were found in localized regions with insufficiently ordered volume to give rise to detectable diffraction spots in the XRD patterns. Similar inter-growth of extra -phases in R-P layers were reported in epitaxial films of SrRuO and SrTiO () grown by MBE [10, 25, 47, 49]. STEM investigations of films grown outside the optimal substrate temperature window showed increased defect density and populations of inter-growths of other -phases of the R-P series [See Fig. S1 of the supplementary material]. These inter-growths are difficult to control in epitaxial growth using PLD. This is because minor stoichiometric deviations that can exist between the target and deposited films, together with the fact that the difference in the formation energies among nearby members of the SrRuO series is small, could lead to the formation of mixed -phase films of these materials [48].
@@
Now we turn to device fabrication and electronic transport characteristics of epitaxial SrRuO films prepared at the optimal growth conditions. To investigate the magnetotransport properties in these films and patterned devices, we fabricated side-by-side on the same film a series of Hall bar devices of various channel widths (from m to nm) [Fig. 5(a)]. The devices were patterned using electron-beam lithography (EBL) followed by ion-beam etching and the deposition of Ti/Au (5 nm/45 nm) electrodes. Figures 5(b) and 5(c) show top-view scanning electron microscopy (SEM) images of patterned Hall bar devices and a schematic side view of a device together with transport measurement configuration, respectively. Unpatterned SrRuO films were measured in a Van der Pauw configuration by wire bonding aluminum wires to the samples’ corners. Electronic transport characteristics of patterned thin-film devices, which were fabricated side-by-side on the same film after transport measurements in a Van der Pauw geometry, were studied in a Hall bar configuration with aluminum wires connected to the electrodes [Fig. 5(d)]. For magnetotransport measurements, we used a Quantum Design PPMS operated at 300–5 K. An excitation current of A was applied. The MR measurements were performed with the magnetic field oriented both along the axis and in the plane of the film.
@@
Figure 6(a) presents the temperature-dependent zero magnetic field sheet resistance () between 5 and 300 K for a 25-nm-thick unpatterned SrRuO film and for two representative patterned Hall bar devices fabricated side-by-side on this same film after Van der Pauw measurements. The unpatterned SrRuO film and devices show metallic behavior over the entire temperature range. At 5 K, the is roughly a factor of 4 higher in unpatterned film than in Hall bar devices. The extracted residual resistivity ratio (RRR), defined as , is RRR for 25 nm-thick films. This value is consistent with reported resistivity data of MBE-grown SrRuO films of similar thickness [10]. From the temperature dependence of the normalized resistances, , at various applied field orientations, the measured temperature dependence is isotropic in the absence of an applied field, while enhanced scattering is observed for in-plane applied field of 5 T [see, Fig. S2 of the supplementary material]. This enhanced scattering near the metamagnetic field transition was observed in a strained SrRuO films [10], and it has been attributed to field-controlled instabilities of the Fermi surface of SrRuO [31].
@@
Figures 6(b)-6(d) show the MR at several temperatures for the same unpatterned film and two Hall bar devices fabricated on the same sample after measurements in Van der Pauw geometry. The magnetoresistance is defined as , where and are the longitudinal resistance measured at a field and at zero field, respectively. MR data were acquired with the field in out-of-plane direction (axis, ), the excitation current being perpendicular to the field (). No hysteresis loop was observed, which indicates that no intrinsic long-range ferromagnetic order occurs. Details on the analysis of MR data are discussed elsewhere [51, 52, 53]
@@
For unpatterned films measured in Van der Pauw geometry at 5 K, a positive MR is measured at low magnetic fields together with a shallow change of slope slightly above 6 T [Fig. 6(b)]. As the temperature is increased further, this shallow slope at high field disappears and MR data show quasi-quadratic field dependence. Positive MR with change of slope from a positive to a negative, resulting into a MR peak around 6 T, has been reported in thin SrRuO single crystal nanosheets measured at 500 mK with axis [30]. The positive MR peak has been associated with the metamagnetic transition in SrRuO [30, 27]. For patterned Hall bar devices, a positive and quasi-quadratic field dependence MR is observed independent of temperature [Fig. 6(c)-6(d)]. Unlike for bulk crystals, we are not able to resolve a MR peak in SrRuO thin films due most probably to impurity in these films, and also because MR data were acquired only down to 5 K and up to 9 T. For SrRuO single crystals, a broad MR peak at 5 K was resolved only for MR data acquired up to 15 T [27].
@@
The MR increases with decreasing temperature and it evolves into quasi-linear behavior in the high field limit. For the unpatterned SrRuO film, a maximum MR of 10% is obtained [Fig. 6(b)]. However, Hall bar devices fabricated on this same film show large positive MR values of 40% and 140% for devices of channel widths of , respectively [Fig. 6(c)-6(d)]. The positive MR for is an indication of paramagnetic behavior of these films, which is consistent with known resistivity increase as magnetic field increases in most materials [54]. While for ferromagnets, MR is negative; for paramagnetic metals, the movement of carriers will be deflected under a Lorentz force, which increases the probability of carrier scattering, thereby increasing the resistance.
@@
The differences revealed by the transport measurement results [Fig. 6(a)] and MR values of unpatterned films and patterned Hall bars [Fig. 6(b)-(d)] is plausibly caused by inhomogeneities in the electron concentration caused by disorder in the sample. This will cause the bulk region to present different types of local current density and different potential landscape, resulting in different evolutions of the longitudinal potential profiles over certain portion of the film. We note that such types of inhomogeneities in disordered conductors is known to give rise to large MR, as described by the Parish and Littlewood model [55, 56] that was used to explain large MR values in a range of materials (e.g., AgTe, AgSe, and BiTe). In this case, the materials were modeled as a network of resistors due to disorder-induced mobility fluctuations [55, 57, 58]. For Hall bars of various channel widths structured on a SrRuO film with some inhomogeneous disordered conduction (e.g., regions with higher conductivity than others), there will be inhomogeneity in the electron concentration over a small area in Hall bars, thus resulting in larger MR values in Hall bar devices as compared to unpatterned SrRuO films. This interpretation is consistent to the observation that the Hall bar device with narrow channel width (D-07 of 5 m width) exhibit large resistance and MR values as compared to the Hall bar device D-03 that has a channel width of 10 m.
@@
For the same film characterized in Figure 6, we have performed angle dependence MR measurements at various temperatures. Figure 7 depicts representative MR data at 5 K for unpatterned SrRuO film. The MR data were acquired for (the angle between the field and the axis of the sample) values of . As increases from , the MR undergoes a gradual transition from positive to negative. For , the MR is negative and it reaches a local minimum at a magnetic field of T, which is consistent with the in-plane metamagnetic transition behavior [10]. MR data of Hall bar device fabricated on the same film exhibit qualitatively similar switching behaviors from positive to negative MR as the magnetic field is applied from out-of-plane (axis) to in-plane direction (plane) [see Fig. S3 of the supplementary material]. A maximum negative MR values of -4% and -70% are obtained for unpatterned film [Fig. 7] and Hall bar device [Fig. S3], respectively.
@@
Supplementary Material
See the supplementary material for additional microstructural characterizations using TEM and magnetotransport measurements of epitaxial SrRuO films.

### 18 Mar 2024 v4

@@ 3. Results and Discussion
The crystalline quality and phase purity of SrRuO films were characterized by XRD. Figure 2(a) shows representative scans for the SrRuO films at various substrate temperatures and laser fluence. For the films grown at optimal growth conditions (TC and laser fluence ), only the substrate peaks and phase-pure family of the film diffraction peaks are resolved, which indicates a high crystallinity and verifies that the SrRuO films were aligned along the c-axis. The Bragg diffraction peaks are well resolved, while the peaks are not present due to the small film thickness and the low structure factors of these peaks [10]. The extracted out-of-plane lattice parameter for epitaxial films grown at the optimal growth conditions is Å. This value is consistent to the bulk axis lattice constant of 20.7194 Å reported previously in single crystals of SrRuO [40, 41, 42, 30]. Extra phases, indicated by in Fig. 2(a), were detected in SrRuO films grown at high substrate temperatures, underlining the role of proper tuning of T for achieving phase stability in these films. These extra phases are associated with the family of the diffraction peaks of the R-P phase [43, 44], in agreement with STEM data (See Fig. S1 in Ref. [45]).
Figure 2(b) shows a closeup view of the scans around the diffraction peak for the SrRuO films grown at various T and laser fluence. For the films grown at the optimal growth conditions, only the peaks are resolved with noticeable thickness fringes, which highlight phase purity and smooth growth. However, we observed shifting of the peaks to lower angles for films that were not grown at the optimal growth conditions. The shifting of the peak demonstrates that the -axis lattice constant expands, indicating inter-growth of other SrRuO phases and change of stability of the SrRuO films. In particular, the SrRuO films grown at T°C exhibited a peak shoulder around the substrate peak, indicating an inter-growth of higher R-P phases, with predominance of the SrRuO phase [47]. These observations are consistent with the STEM data (see, Fig. S1 in Ref. [45]) in which considerable inter-growths of other members of the SrRuO series were clearly detected for epitaxial SrRuO films prepared outside the optimal growth conditions.
@@
For epitaxial SrRuO films grown at optimal growth conditions (TC and laser fluence ), the defects detected are infrequently regions of inter-growth within layers (indicated by black arrows in Figs. 3(a) and 3(c)). Such inter-growths were found in localized regions with insufficiently ordered volume to give rise to detectable diffraction spots in the XRD patterns. Similar inter-growth of extra -phases in R-P layers were reported in epitaxial films of SrRuO and SrTiO () grown by MBE [10, 25, 48, 50]. STEM investigations of films grown outside the optimal substrate temperature window showed increased defect density and populations of inter-growths of other -phases of the R-P series (See Fig. S1 of Ref. [45]). These inter-growths are difficult to control in epitaxial growth using PLD. This is because minor stoichiometric deviations that can exist between the target and deposited films, together with the fact that the difference in the formation energies among nearby members of the SrRuO series is small, could lead to the formation of mixed -phase films of these materials [49].
@@
Figure 5(a) presents the temperature-dependent zero magnetic field sheet resistance () between 5 and 300 K for a 25-nm-thick unpatterned SrRuO film and for two representative patterned Hall bar devices fabricated side-by-side on this same film after Van der Pauw measurements. (For details on Hall bar devices, see Fig. S3 of Ref. [45]). The unpatterned SrRuO film and devices show metallic behavior over the entire temperature range. At 5 K, the is roughly a factor of 4 higher in unpatterned film than in Hall bar devices. The extracted residual resistivity ratio (RRR), defined as , is RRR for 25 nm-thick films. This value is consistent with reported resistivity data of MBE-grown SrRuO films of similar thickness [10]. From the temperature dependence of the normalized resistances, , at various applied field orientations, the measured temperature dependence is isotropic in the absence of an applied field, while enhanced scattering is observed for in-plane applied field of 5 T (see, Fig. S2 of Ref. [45]). This enhanced scattering near the metamagnetic field transition was observed in a strained SrRuO films [10], and it has been attributed to field-controlled instabilities of the Fermi surface of SrRuO [31].
@@
Figures 5(b)-5(d) show the MR at several temperatures for the same unpatterned film and two Hall bar devices fabricated on the same sample after measurements in Van der Pauw geometry. The magnetoresistance is defined as , where and are the longitudinal resistance measured at a field and at zero field, respectively. MR data were acquired with the field in out-of-plane direction (axis, ), the excitation current being perpendicular to the field (). No hysteresis loop was observed, which indicates that no intrinsic long-range ferromagnetic order occurs. Details on the analysis of MR data are discussed elsewhere [52, 53, 54]
@@
For unpatterned films measured in Van der Pauw geometry at 5 K, a positive MR is measured at low magnetic fields together with a shallow change of slope slightly above 6 T [Fig. 5(b)]. As the temperature is increased further, this shallow slope at high field disappears and MR data show quasi-quadratic field dependence. Positive MR with change of slope from a positive to a negative, resulting into a MR peak around 6 T, has been reported in thin SrRuO single crystal nanosheets measured at 500 mK with axis [30]. The positive MR peak has been associated with the metamagnetic transition in SrRuO [30, 27]. For patterned Hall bar devices, a positive and quasi-quadratic field dependence MR is observed independent of temperature [Fig. 5(c)-5(d)]. Unlike for bulk crystals, we are not able to resolve a MR peak in SrRuO thin films due most probably to impurity in these films, and also because MR data were acquired only down to 5 K and up to 9 T. For SrRuO single crystals, a broad MR peak at 5 K was resolved only for MR data acquired up to 15 T [27].
@@
The MR increases with decreasing temperature and it evolves into quasi-linear behavior in the high field limit. For the unpatterned SrRuO film, a maximum MR of 10% is obtained [Fig. 5(b)]. However, Hall bar devices fabricated on this same film show large positive MR values of 40% and 140% for devices of channel widths of , respectively [Fig. 5(c)-5(d)]. The positive MR for is an indication of paramagnetic behavior of these films, which is consistent with known resistivity increase as magnetic field increases in most materials [55]. While for ferromagnets, MR is negative; for paramagnetic metals, the movement of carriers will be deflected under a Lorentz force, which increases the probability of carrier scattering, thereby increasing the resistance.
@@
The differences revealed by the transport measurement results [Fig. 5(a)] and MR values of unpatterned films and patterned Hall bars [Fig. 5(b)-(d)] is plausibly caused by inhomogeneities in the electron concentration caused by disorder in the sample. This will cause the bulk region to present different types of local current density and different potential landscape, resulting in different evolutions of the longitudinal potential profiles over certain portion of the film. We note that such types of inhomogeneities in disordered conductors is known to give rise to large MR, as described by the Parish and Littlewood model [56, 57] that was used to explain large MR values in a range of materials (e.g., AgTe, AgSe, and BiTe). In this case, the materials were modeled as a network of resistors due to disorder-induced mobility fluctuations [56, 58, 59]. For Hall bars of various channel widths structured on a SrRuO film with some inhomogeneous disordered conduction (e.g., regions with higher conductivity than others), there will be inhomogeneity in the electron concentration over a small area in Hall bars, thus resulting in larger MR values in Hall bar devices as compared to unpatterned SrRuO films. This interpretation is consistent to the observation that the Hall bar device with narrow channel width (D-07 of 5 µm width) exhibit large resistance and MR values as compared to the Hall bar device D-03 that has a channel width of 10 µm.
@@
For the same film characterized in Figure 5, we have performed angle dependence MR measurements at various temperatures. Figure 6 depicts representative MR data at 5 K for unpatterned SrRuO film. The MR data were acquired for (the angle between the field and the axis of the sample) values of . As increases from , the MR undergoes a gradual transition from positive to negative. For , the MR is negative and it reaches a local minimum at a magnetic field of T, which is consistent with the in-plane metamagnetic transition behavior [10]. MR data of Hall bar device fabricated on the same film exhibit qualitatively similar switching behaviors from positive to negative MR as the magnetic field is applied from out-of-plane (axis) to in-plane direction (plane) (see Fig. S4 of Ref.[45]). A maximum negative MR values of -4% and -70% are obtained for unpatterned film [Fig. 6] and Hall bar device (see Fig. S4 of Ref.[45]), respectively.